OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4.

OpenStudy (anonymous):

Which of the following is an example of negative correlation? Time worked vs money earned House address vs math test score Age vs commute time to school Time eating vs food left

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (oaktree):

Phew! OK, for the first one, you should use the same slope (-3) and the point given and plug it into y=mx+b to solve for b.

OpenStudy (oaktree):

For the second, reason it out. The more time you work, what happens to the money you earn? Is there even a relationship between the second and the third ones? The more you eat, what happens to the amount of food you have left?

OpenStudy (oaktree):

For the last one, just plug in the points to each equation and check which one comes out to a correct statement.

OpenStudy (anonymous):

what to do mean